OpenStudy (chosenmatt):

what is the difference? 1/9-5/6

OpenStudy (anonymous):

1/9 - 5/6 take lcm of 9, 6 = 18 2-5(3)/18= 2-15/18 ans = -13/18

OpenStudy (anonymous):



OpenStudy (deadpoolroxz):

dey stright